A capacitor consists of two short sections of parallel wires in air. The voltage on one wire is 60 V and on the other is -60 V. The two wires carry equal and opposite charge with the magnitude of 476 pC. Find the capacitance of the capacitor in pF
Answer #1

capacitance is C = Q/V

Q is the charge on the plates = 476 pC


V is the voltage across plates = 60 V


then Capacitance is C = (476*10^-12)/60 = 7.93*10^-12 F = 7.93 pF